Hi need some help with my daughters homework. Here is the question:
Identify the slope for each line of best fit.
y= 3x + 4
y= 2X - 8
y= -4 x +6
y= 11x
y= .35x - 7
Any help would be wonderful. Thanks Chuck

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

Identify the slope for each line of best fit.
y= 3x + 4
y= 2X - 8
y= -4 x +6
y= 11x
y= .35x - 7
--------------
The slope-intercept form of equations is y = mx+b, where m is the slope.
These 5 are in that form, so the slopes are the coefficients of x.
y= 3x + 4 m = 3
y= 2X - 8 m = 2
y= -4 x +6 m = -4
y= 11x m = 11
y= .35x - 7 m = .35




Answer by checkley77(12844)

y= 3x + 4 slope=3
y= 2X - 8 slope=2
y= -4x +6 slope=-4
y= 11x slope=11
y= .35x - 7 slope=.35